# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v3

@@ 3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.5)
@@
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the unbiased sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h(h-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h}{h-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.6)
where $\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean of the observed edges. Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 4 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$. The key advantage of this approach is that it completely bypasses the need for asymptotic plug-iin justifications, providing a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ based entirely on the observable quantity $\hat{\sigma}_{ij}^{2}$.
@@
For any $t>0$,
$\mathbb{P}\Bigg(\big|U_{ij}^{(-j)}\big|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\,t}{h}}+\frac{7t}{3(h-1)}\ \Bigg|\ \mathcal{N}_{i}^{(-j)},P\Bigg)\leq 2e^{-t}.$
@@
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
and its conditional expectation is
$\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}$ is exactly the deviation of the empirical mean from its conditional expectation. Applying Theorem 4 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, and taking a union bound over both tails yields the stated inequality. ∎
Lemma 3.8 provides a strictly finite-sample, variance-adaptive control of $U_{ij}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby completely bypassing the need for asymptotic plug-in justifications.
@@ 3.3 Berry-Esseen normal approximation
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
@@
Suppose Assumptions 2.1 and 2.2 hold and $h_{n}\asymp\sqrt{n\log n}$. Then there exists a universal constant $C>0$, independent of $i,j,n$, such that, uniformly over $1\leq i,j\leq n$,
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.9)
@@
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Feller (1971, Chapter XVI)) states that there exists a universal constant $C_{0}>0$ such that
@@
By (3.8), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$ and therefore it does not depend on $i$ and $j$. Finally, since
$U_{ij}^{(-j)}=\frac{1}{h_{n}}S,\qquad V_{ij}=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2},$
we have the identity
@@
and the bound (3.9) follows. ∎
Corollary 3.10 (Asymptotic normality).
@@
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,1)\qquad\text{uniformly over }1\leq i,j\leq n.$
@@
Immediate from Theorem 3.9, since $h_{n}^{-1/2}\to 0$. ∎
@@ 3.4 Confidence intervals
Undder the canonical choice $h_{n}\asymp\sqrt{n\log n},$ we have
@@
By Lemma 3.7, the bias component satisfies
$\sup_{i,j}|B_{ij}^{(-j)}|=O\!\big((\log n/n)^{1/4}\big).$
Hence the stochastic fluctuation scale and the bias scale are of the same order, up to logarithmic factors. In particular, to obtain asymptotically valid normal-approximation confidence intervals for $P_{ij}$, one must either undersmooth (so that the bias is $o(h_{n}^{-1/2})$) or incorporate an explicit bias correction of order $(\log n/n)^{1/4}$.
@@ 3.4.2 Normal‑approximation intervals
By Theorem 3.9 and the bias control, the coverage of $\mathrm{CI}_{ij}^{(N)}(\alpha)$ differs from $1-\alpha$ by at most $O\big((n\log n)^{-1/4}\big)$, up to the error from estimating $V_{ij}$.
@@
1:Sample size $n$, neighborhood size $h$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
@@
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=h(\xi_{i},\xi_{j})$
@@ 4 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h)$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h)$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard
@@ 4.1 Cross-validation procedure and notation
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h):=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h)}A_{kj},\qquad j\neq i.$
The empirical cross-validation score at row $i$ is the average squared prediction error across all available held-out columns:
$\widehat{R}_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}.$ (4.1)
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h)$ given $P$, this evaluates to
$R_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h),P\right].$ (4.2)

### v4

@@ 3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.9)
@@
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the (Bessel-corrected) sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.10)
where $\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean.
Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 11 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$.
For the normal-approximation intervals we estimate the conditional variance
$V_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj})$
by the plug-in estimator
$\widehat{V}_{ij}:=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$
which replaces the unknown probabilities $P_{kj}$ by their smoothed estimates.
@@
Assume $h_{n}\geq 2$. For any $\delta\in(0,1)$,
$\mathbb{P}\!\left(|U_{ij}^{(-j)}|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\ln(2/\delta)}{h_{n}}}+\frac{7\ln(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\leq\delta.$
@@
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Moreover, by the leave-one-out construction, $\mathcal{N}_{i}^{(-j)}$ depends only on $A^{(-j)}$ and is independent of $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ given $P$.
Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
with conditional expectation
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}^{(-j)}$ is exactly the deviation of the empirical mean from its conditional expectation.
Applying Theorem 11 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$ and again to $1-X_{k}$ yields the stated two-sided inequality. ∎
Lemma 3.13 provides a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby bypassing the need for asymptotic plug-in arguments.
@@ 3.3 Berry-Esseen normal approximation
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid\mathcal{N}_{i}^{(-j)},P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
@@
Lemma 3.14 (Uniform plug-in control for $\widehat{V}_{ij}$).
Recall
$V_{ij}=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),\qquad\widehat{V}_{ij}=\frac{1}{h_{n}^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}).$
Then for all $i\neq j$,
$\big|\widehat{V}_{ij}-V_{ij}\big|\leq\frac{1}{h_{n}}\,\sup_{1\leq a,b\leq n}\big|\widehat{P}_{ab}-P_{ab}\big|.$
Consequently, if one had uniform sup-norm consistency
$\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|=o_{p}(1),$
then the plug-in variance estimator $\widehat{V}_{ij}$ would inherit ratio-consistency. However, we do not establish such a sup-norm bound in this paper. Rowwise mean-squared error control does not imply uniform entrywise consistency. Therefore we do not claim ratio-consistency of $\widehat{V}_{ij}$, and normal intervals using $\widehat{V}_{ij}$ incur an additional, unquantified variance-estimation error.
Proof.
The map $\psi(x)=x(1-x)$ is $1$-Lipschitz on $[0,1]$ since $|\psi^{\prime}(x)|=|1-2x|\leq 1$. Hence for any $k$,
$\big|\widehat{P}_{kj}(1-\widehat{P}_{kj})-P_{kj}(1-P_{kj})\big|\leq|\widehat{P}_{kj}-P_{kj}|\leq\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|.$
Summing over $k\in\mathcal{N}_{i}^{(-j)}$ and dividing by $h_{n}^{2}$ gives
$|\widehat{V}_{ij}-V_{ij}|\leq\frac{1}{h_{n}^{2}}\cdot h_{n}\cdot\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|=\frac{1}{h_{n}}\sup_{a,b}|\widehat{P}_{ab}-P_{ab}|.$
If $V_{ij}\asymp h_{n}^{-1}$ uniformly (e.g., under Assumption 2.2), then dividing by $V_{ij}$ yields the ratio consistency claim. ∎
Remark 3.15.
Our theoretical results establish rowwise mean-squared error control but do not prove uniform sup-norm consistency of $\widehat{P}$. Accordingly, plug-in variance intervals based on $\widehat{V}_{ij}$ should be interpreted as heuristic refinements rather than formally guaranteed procedures.
To control the additional error induced by replacing $V_{ij}$ by $\widehat{V}_{ij}$, we use Lemma 3.14, which reduces variance plug-in accuracy to the uniform estimation error of $\widehat{P}$.
@@
Suppose Assumptions 2.1 and 2.2 hold, and let $h_{n}\to\infty$ as $n\to\infty$. Then there exists a universal constant $C>0$, independent of $i,j,n$, such that
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.13)
@@
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Petrov (1975, Chapter V, Theorem 3)) states that there exists a universal constant $C_{0}>0$ such that
@@
By (3.12), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$, where $C$ is independent of $(i,j)$. Finally, since $U_{ij}^{(-j)}=h_{n}^{-1}S$ and $V_{ij}=h_{n}^{-2}\sigma_{ij}^{2}$, we have the identity
@@
and the uniform bound (3.13) follows. ∎
Corollary 3.17 (Uniform asymptotic normality).
@@
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\longrightarrow 0\qquad\text{as }n\to\infty.$
@@
For fixed $(i,j)$ and $x\in\mathbb{R}$, define
$\Delta_{ij}(x):=\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x).$
By Theorem 3.16,
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}|\Delta_{ij}(x)|\leq Ch_{n}^{-1/2}\qquad\text{a.s.}$
Taking expectations and using Jensen’s inequality for $|\,\cdot\,|$,
$\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|=\left|\mathbb{E}[\Delta_{ij}(x)]\right|\leq\mathbb{E}[|\Delta_{ij}(x)|]\leq Ch_{n}^{-1/2}.$
Taking the supremum over $x\in\mathbb{R}$ and $(i,j)$ yields
$\sup_{1\leq i,j\leq n}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2},$
and since $h_{n}\to\infty$, the right-hand side tends to 0. ∎
@@ 3.4 Confidence intervals
Under the canonical choice $h_{n}\asymp\sqrt{n\log n},$ we have
@@
By Lemma 3.11, the bias component satisfies Consequently, Corollary 3.9 yields a uniform bias bound over $(i,j)\in\mathcal{E}_{\varepsilon}$. Hence the stochastic fluctuation scale and the bias scale are of the same order, up to logarithmic factors. In particular, to obtain asymptotically valid normal-approximation confidence intervals for $P_{ij}$, one must either undersmooth (so that the bias is $o(h_{n}^{-1/2})$) or incorporate an explicit bias correction of order $(\log n/n)^{1/4}$.
@@ 3.4.1 Finite-sample Confidence Intervals via Empirical Bernstein Bounds
Theorem 3.18 (Finite-sample inference for $P_{ij}$ on the good set).
Fix $\varepsilon>0$, $\alpha\in(0,1)$, and $(i,j)\in\mathcal{E}_{\varepsilon}$. Let $CI^{(B)}_{ij}(\alpha)$ be the empirical Bernstein interval defined in (3.17) (using the sample variance form). On the event that the neighborhood regularity condition (3.4) holds (probability at least $1-n^{-c}$ as in Corollary 3.9), we have the bias cushion
$\left|P_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\right|\leq C(\varepsilon,C_{2})\,r_{n}.$
Moreover, by the empirical Bernstein deviation (Maurer and Pontil (2009)) in (3.17)
$\mathbb{P}\!\left(\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\in CI^{(B)}_{ij}(\alpha)\right)\geq 1-\alpha.$
Hence, on the intersection of these two events,
$P_{ij}\in CI^{(B)}_{ij}(\alpha)\oplus[-C(\varepsilon,C_{2})r_{n},\;C(\varepsilon,C_{2})r_{n}].$
Consequently,
$\mathbb{P}\!\big(P_{ij}\in CI^{(B)}_{ij}(\alpha)\oplus[-C(\varepsilon,C_{2})r_{n},\;C(\varepsilon,C_{2})r_{n}]\big)\geq 1-\alpha-n^{-c}.$
Remark 3.19 (Target of the Empirical Bernstein interval).
It is important to note that $CI_{ij}^{(B)}(\alpha)$ is centered at the empirical neighborhood mean $\tilde{P}_{ij}$ and therefore provides finite-sample coverage for the localized average
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj},$
not for the individual probability $P_{ij}$. The distinction reflects the fact that the Empirical Bernstein inequality controls only the stochastic fluctuation component $U_{ij}^{(-j)}$ and does not address the bias term $B_{ij}^{(-j)}$.
@@ 3.4.2 Normal‑approximation intervals
By Theorem 3.16, the uniform Berry-Esseen bound implies that
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\left|\mathbb{P}\!\big(P_{ij}\in\mathrm{CI}_{ij}^{(N)}(\alpha)\big)-(1-\alpha)\right|=O\!\big(h_{n}^{-1/2}\big).$
Remark 3.21.
The Berry-Esseen bound for the variance component holds uniformly over all $(i,j)$. However, coverage statements for $P_{ij}$ require bias control and therefore hold uniformly only over $(i,j)\in\mathcal{E}_{\varepsilon}$.
These intervals are asymptotically valid for the variance component $U_{ij}^{(-j)}$ uniformly in $(i,j)$, and for $P_{ij}$ uniformly over $(i,j)\in\mathcal{E}_{\varepsilon}$ up to the bias cushion. Under the canonical choice $h_{n}\asymp\sqrt{n\log n}$ this becomes $O\!\big((n\log n)^{-1/4}\big)$. The bias cushion in (3.20) ensures that the bias term from Lemma 3.11 is asymptotically negligible at the same scale. Normal intervals using $\widehat{V}_{ij}$ therefore incur an additional variance-estimation error term, as discussed in Lemma 3.14.
@@
1:Sample size $n$, neighborhood size $h_{n}$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
@@
3:Draw $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ and set $P_{ij}=f(\xi_{i},\xi_{j})$
@@ 4 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h_{n}$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h_{n})$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard concentration and normal approximation results, enabling valid CV-based tuning of $h_{n}$.
We use the term “honest” to denote that no held-out edge $A_{ij}$ is used in constructing its predictor $\tilde{P}_{ij}^{(-j)}$. This eliminates double use of data but does not imply oracle optimality of the selected bandwidth.
@@ 4.1 Cross-validation procedure and notation
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h_{n}\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h_{n})$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h_{n}):=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h_{n})}A_{kj},\qquad j\neq i.$
Define the LOO cross-validation score
$\widehat{R}_{i}(h)=\frac{1}{n-1}\sum_{j\neq i}(A_{ij}-\tilde{P}_{ij}^{(-j)}(h))^{2}.$
Because the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed from $A^{(-j)}$ (which omits all edges incident to vertex $j$) and $i\notin\mathcal{N}_{i}^{(-j)}(h)$ by construction, the held‑out edge $A_{ij}$ is conditionally independent of the predictor $\tilde{P}_{ij}^{(-j)}(h)$ given $P$. Consequently, $\widehat{R}_{i}(h)$ is (conditionally) unbiased for the corresponding prediction risk. A full oracle inequality analysis of this tuning rule is not included in this version.
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h_{n})$ given $P$, this evaluates to
$R_{i}(h_{n}):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h_{n}),P\right].$ (4.1)